OpenStudy (anonymous):

A small child makes 3 piles of blocks in a room. Each pile has 4 blocks. In another room, he makes 6 more piles of blocks. These piles also have 4 blocks each. To count the total number of blocks, his sister counts that there are 3 times 4, or 12 blocks, in one room and 6 times 4, or 24 blocks, in the other room for a total of 36 blocks. Using the distributive property, what is another way she could have counted the blocks?

OpenStudy (anonymous):

3x4 + 6x4 = (3+6) x 4 = 36

OpenStudy (anonymous):

3 times 4 + 6 times 4 = (3+6) times 4 = 9 times 4 = 36

OpenStudy (anonymous):

is it okay?

OpenStudy (anonymous):

@OneKoi ...?

OpenStudy (anonymous):

It's okay. I forgot to add the options, though: a) She could have counted each block separately. b) She could have switched the order in which she added the 12 and 24. c) She could have found the number of blocks in 4 piles and added it to the number of blocks in 5 piles. d) She could have added the number of piles before multiplying by 4.

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Thanks.